# CSI-BERT2: A BERT-inspired Framework for Efficient CSI Prediction and Classification in Wireless Communication and Sensing

### v2

@@ Abstract
Channel state information (CSI) is a fundamental component in both wireless communication and sensing systems, enabling critical functions such as radio resource optimization and environmental perception. In wireless sensing, data scarcity and packet loss hinder efficient model training, while in wireless communication, high-dimensional CSI matrices and short coherent times caused by high mobility present challenges in CSI estimation. To address these issues, we propose a unified framework named CSI-BERT2 for CSI prediction and classification tasks. Building on CSI-BERT, we introduce a two-stage training method that first uses a mask language model (MLM) to enable the model to learn general feature extraction from scarce datasets in an unsupervised manner, followed by fine-tuning for specific downstream tasks. Specifically, we extend MLM into a mask prediction model (MPM), which efficiently addresses the CSI prediction task. We also introduce an adaptive re-weighting layer (ARL) to enhance subcarrier representation and a multi-layer perceptron (MLP)-based temporal embedding module to mitigate permutation invariance issues in time-series CSI data. This significantly improves the CSI classification performance of the original CSI-BERT model. Extensive experiments on both real-world collected and simulated datasets demonstrate that CSI-BERT2 achieves state-of-the-art performance across all tasks. Our results further show that CSI-BERT2 generalizes effectively across varying sampling rates and robustly handles discontinuous CSI sequences caused by packet loss—challenges that conventional methods fail to address.
Index Terms:
Channel statement information (CSI), CSI prediction, CSI classification, wireless communication, wireless sensing
@@ I Introduction
Channel state information (CSI) plays a critical role in both wireless sensing and wireless communication systems by capturing the propagation characteristics of the wireless channel. This information provides valuable insights into the radio environment, including the channel conditions, signal quality, and the motion and location of objects. In wireless communication systems, CSI is essential for optimizing key processes in wireless communication systems, such as channel compensation, adaptive modulation and coding, user selection and scheduling. Effective utilization of CSI enables improved system performance, increased spectrum efficiency, and more reliable communication, ultimately enhancing the overall user experience [1, 2, 3]. In wireless sensing applications, CSI can enhance the ability to detect and track objects, leading to improved environmental awareness. Owing to its advantages in privacy protection, low cost, and penetration ability, wireless sensing has been widely adopted in diverse applications, such as fall detection [4], localization [5], and people identification [6].
@@
To address the aforementioned shortcomings, we propose a method inspired by bidirectional encoder representations from transformers (BERT) [18], named CSI-BERT2, which is an improved version of our previous work, CSI-BERT [7]. In CSI-BERT, the masked language model (MLM) approach was first introduced for the pre-training of the CSI recovery model in Wi-Fi sensing, enabling it to learn effectively from limited dataset and improving performance in downstream tasks. Building on this, CSI-BERT2 introduces the mask prediction model (MPM), where we mask the final portion of the CSI sequence and train the model to recover it. By predicting the masked CSI in a single step, our method efficiently avoids the error accumulation problem common in AR models. Additionally, we propose a novel multi-layer perceptron (MLP)-based time embedding method, effectively addressing the permutation invariance problem [19]. Unlinke traditional positional embeddings in Transformers that captures only distance relationships, our proposed time embedding method explicitly exploit the temporal order, which is critical for CSI sequences. This design enables CSI-BERT2 to generalize efficiently in environments with varying sampling rates, and perform robustly even in a zero-shot setting. Finally, considering the varying information content convened across different subcarriers, we introduce an adaptive re-weighting layer (ARL) [20], which helps to adaptively capture the importance of each subcarrier. Through these architectural enhancements, we successfully address the low performance of the original CSI-BERT in CSI classification tasks. In conclusion, the main contributions of this paper are summarized as follows:
@@
(2) Improved Model Structure: We design a novel time embedding technique that utilizes MLP to further encode temporal information, addressing the permutation invariance problem inherent in the Transformer positional embedding layer. This design enables our model to function effectively in various sampling rate scenarios. Furthermore, we propose a simple attention mechanism called ARL to better capture the relationships and importance between subcarriers adaptively.
@@
(4) Experiment Evaluation: To validate the performance of CSI-BERT2, we conduct comprehensive experiments across four datasets, including both public and self-collected data from real-world and simulated environments. In the CSI prediction task, CSI-BERT2 achieves the lowest error, outperforming the current state-of-the-art (SOTA) by over 60% in mean squared error (MSE). Although it requires slightly more computation time than other models, the total prediction time for the entire dataset, consisting of hundreds of samples, remains under 0.5 seconds, demonstrating its practicality for real-time deployment. In the CSI classification task, CSI-BERT2 consistently achieves the highest accuracy across various applications, including gesture recognition, people identification, fall detection, and crowd counting. Notably, our method shows strong robustness to discontinuous data caused by packet loss and to sampling rate mismatches between the training and testing sets, without the need to process and recover data or train a new model separately. Its accuracy drops by no more than 2%, in contrast to other models that suffer tens of percentage points in degradation. By demonstrating robust performance under packet loss and variable sampling rates, CSI-BERT2 eliminates the need for resource-intensive retraining in dynamic environments, making it viable for real-time applications such as emergency fall detection.
@@ III-A Overview
However, the original CSI-BERT does not fully leverage the advantages of MLM, resulting in subpar performance in downstream tasks. For example, the classification accuracy of CSI-BERT after fine-tuning does not surpass that of the vanilla ResNet18 [36]. This limitation can be attributed to the inadequacies of the embedding layer. To address this issue, CSI-BERT2 introduces an ARL-based spatial encoding module to capture the relationships and importance among subcarriers. Additionally, a MLP-based temporal encoding module is proposed to capture both the order and distance relationships among packets, effectively solving the permutation invariance problem inherent in conventional positional embedding modules, which capture distance relationships without considering order information. Moreover, to fully utilize the learned knowledge during pre-training, we propose a novel fine-tuning method named mask prediction model (MPM) for the CSI prediction task, wherein the last portion of the CSI sequence is randomly masked, and the model is trained to predict these masked values. Compared to conventional AR-based methods, our approach can complete the prediction in one step, thereby eliminating the error accumulation problem.
The model structure and working scenarios of CSI-BERT2 are illustrated in Fig. 1 and Fig. 2, respectively. We leverage the CSI sequence $c=[c_{1},c_{2},\ldots,c_{N}]$ and the corresponding timestamps $t=[t_{1},t_{2},\ldots,t_{N}]$ as input. Each $c_{i}$ represents a flattened vector of the CSI matrix at time $t_{i}$, which contains the amplitude and phase information of the subcarriers.
@@ III-B Data Pre-processing
Due to packet loss, the received CSI series are always incomplete. Before inputting them to our model, we need to first identify where the packet loss happens and use a placeholder [PAD] to occupy the place of lost packets. We determine the occurrence of packet loss by analyzing the time gap between two successive timestamps. Given a sampling rate $f$, the ideal time gap between two successive timestamps should be $\Delta t=\frac{1}{f}$ seconds. Assume we receive a CSI series $[c_{1},c_{2},\dotsc,c_{n}]$ with corresponding timestamps $[t_{1},t_{2},\dotsc,t_{n}]$ with the unit of seconds. Then we can identify whether there is packet loss and how many packets were lost between $c_{i}$ and $c_{i+1}$. We calculate the lost packet amount according to:
@@ III-C1 Spatial and Temporal Embedding Layer
Compared to the spatial encoding, which primarily focuses on the inner information within each packet, the temporal encoding aims to provide more order and position information to the BERT backbone for better learning of the relationships among packets. However, we noted that the vanilla positional embedding in Transformers is insufficient. First, the arrival gap between successive packets is not uniform due to various factors such as environmental noise and movement. The positional embedding cannot capture these small differences. Most importantly, it has been shown to be weak in time series data [19]. Due to the permutation-invariant nature of the Transformer architecture, it can only capture distance information but not order information.
@@ III-C2 Dual-branch Head Layer
For the sequence-level head, we use the same architecture as most Transformer-based methods [39], which first incorporates a self-attention layer to combine the features in the time dimension and then uses a linear layer to generate the final output.
@@ III-D Unsupervised Pre-training Process
For CSI sequence, the MLM involves filling empty positions with [PAD] tokens, randomly masking tokens with [MASK] tokens, and training the model to recover the masked tokens. In BERT, the value of [MASK] token is fixed, but in CSI-BERT2, we replace the word embedding layer with a linear layer, which means the [MASK] token can take on different values affecting the output. To address this, we assign a random value to [MASK] tokens, sampled from a Gaussian distribution $[MASK]_{i}^{(j)}\sim\text{N}(\mu_{i}^{(j)},\sigma_{i}^{(j)})$, where the mean and standard deviation are calculated from the standardization layer. This random approach makes it more challenging for the model to identify the positions of the [MASK] tokens, thereby encouraging a deeper understanding of the CSI sequence structure. For [PAD] token, even though we use a fixed value for [PAD], we replace them with [MASK] tokens before feeding the CSI sequence to the model. This is because there is no [PAD] in the CSI recovery phase. Even though the [PAD] can be ignored by the attention mechanism through the attention mask matrix, it still affects the output of the ARL, which works in the dimension of subcarrier. If [PAD] tokens only appear in the unsupervised training phase, it would cause a gap between inference and training, thereby influencing the model’s performance. As a result, we also disable the attention matrix mechanism in this phase.
@@
where $D$ and $R$ represent the discriminator and recoverer, respectively, and $c$ represents the original CSI input. The discriminator tries to distinguish the original CSI input and the recovered CSI sequence output by the recoverer. The recoverer, in turn, learns to generate outputs that are indistinguishable from real CSI.
$L_{1}=\text{MSE}(c,\hat{c})\ ,$ (9)
$L_{2}=\text{MSE}(\mu,\hat{\mu})\ ,$
$L_{3}=\text{MSE}(\sigma,\hat{\sigma})\ ,$
@@
To reduce overfitting, we also modify the training data selection strategy. Unlike CSI-BERT, which uses a sliding window, CSI-BERT2 randomly samples CSI sequences from the entire dataset during training.
@@ III-E1 CSI Prediction
In the inference phase of the prediction task, we first use [MASK] to replace all [PAD] tokens, and then we fix the last 20% of tokens in the input sequences as [MASK]. The model is then used to predict the last 20% of tokens.
@@ III-E2 CSI Classification
For CSI classification tasks, we adopt the approach commonly used for pre-trained models, where we freeze the bottom layers of CSI-BERT2 and only train the top layers for task-specific adaptation. Different from CSI-BERT, in this phase, we use [MASK] tokens to replace not only all [PAD] tokens but also some other input tokens randomly, to imitate the situation of different packet loss rates. This improves generalization to real-world scenarios where packet loss rates may vary during deployment.

### v3

@@ Abstract
Channel state information (CSI) is a fundamental component in both wireless communication and sensing systems, enabling critical functions such as radio resource optimization and environmental perception. In wireless sensing, data scarcity and packet loss hinder efficient model training, while in wireless communication, high-dimensional CSI matrices and short coherent times caused by high mobility present challenges in CSI estimation. To address these issues, we propose a unified framework named CSI-BERT2 for CSI prediction and classification tasks, built on CSI-BERT, which adapts BERT to capture the complex relationships among CSI sequences through a bidirectional self-attention mechanism. We introduce a two-stage training method that first uses a mask language model (MLM) to enable the model to learn general feature extraction from scarce datasets in an unsupervised manner, followed by fine-tuning for specific downstream tasks. Specifically, we extend MLM into a mask prediction model (MPM), which efficiently addresses the CSI prediction task. To further enhance the representation capacity of CSI data, we modify the structure of the original CSI-BERT. We introduce an adaptive re-weighting layer (ARL) to enhance subcarrier representation and a multi-layer perceptron (MLP)-based temporal embedding module to mitigate temporal information loss problem inherent in the original Transformer. Extensive experiments on both real-world collected and simulated datasets demonstrate that CSI-BERT2 achieves state-of-the-art performance across all tasks. Our results further show that CSI-BERT2 generalizes effectively across varying sampling rates and robustly handles discontinuous CSI sequences caused by packet loss—challenges that conventional methods fail to address. The dataset and code are publicly available at https://github.com/RS2002/CSI-BERT2.
@@ I Introduction
Channel state information (CSI) plays a critical role in both wireless sensing and wireless communication systems by capturing the propagation characteristics of the wireless channel. This information provides valuable insights into the radio environment, including the channel conditions, signal quality, and the motion and location of objects. In wireless communication systems, CSI is essential for optimizing key processes in wireless communication systems, such as channel compensation, adaptive modulation and coding, user selection and scheduling. Effective utilization of CSI enables improved system performance, increased spectrum efficiency, and more reliable communication, ultimately enhancing the overall user experience [1, 2, 3]. In wireless sensing applications, CSI can enhance the ability to detect and track objects, leading to improved environmental awareness [4, 5, 6]. Owing to its advantages in privacy protection, low cost, and penetration ability, wireless sensing has been widely adopted in diverse applications, such as fall detection [7], localization [8], gesture recognition [9], and person identification [10].
@@
To address the aforementioned shortcomings, we propose a method inspired by bidirectional encoder representations from transformers (BERT) [22], named CSI-BERT2, which is an improved version of our previous work, CSI-BERT [11]. In CSI-BERT, the masked language model (MLM) approach was first introduced for the pre-training of the CSI recovery model in Wi-Fi sensing, enabling it to learn effectively from limited dataset and improving performance in downstream tasks. Building on this, CSI-BERT2 introduces the mask prediction model (MPM), where we mask the final portion of the CSI sequence and train the model to recover it. By predicting the masked CSI in a single step, our method efficiently avoids the error accumulation problem common in AR models. Additionally, we propose a novel multi-layer perceptron (MLP)-based time embedding method that effectively addresses the temporal information loss problem [23] inherent in the original Transformer. As illustrated in [23], the permutation-invariant nature of the attention mechanism leads to temporal information loss, even with positional embeddings. In contrast, our proposed time embedding method explicitly exploits the temporal order, which is critical for CSI sequences, thereby mitigating this issue. This design enables CSI-BERT2 to generalize efficiently in environments with varying sampling rates, and perform robustly even in a zero-shot setting. Finally, considering the varying information content convened across different subcarriers, we introduce an adaptive re-weighting layer (ARL) [24], which helps to adaptively capture the importance of each subcarrier. Through these architectural enhancements, we successfully address the low performance of the original CSI-BERT in CSI classification tasks. In conclusion, the main contributions of this paper are summarized as follows:
@@
(2) Improved Model Structure: We design a novel time embedding technique that utilizes MLP to further encode temporal information, addressing the temporal information loss problem inherent in the Transformer positional embedding layer. This design enables our model to function effectively in various sampling rate scenarios. Furthermore, we propose a simple attention mechanism called ARL to better capture the relationships and importance between subcarriers adaptively.
@@
(4) Experiment Evaluation: To validate the performance of CSI-BERT2, we conduct comprehensive experiments across four datasets, including both public and self-collected data from real-world and simulated environments. In the CSI prediction task, CSI-BERT2 achieves the lowest error, outperforming the current state-of-the-art (SOTA) by over 60% in mean squared error (MSE). Although it requires slightly more computation time than other models, the total prediction time for the entire dataset, consisting of hundreds of samples, remains under 0.5 seconds, demonstrating its practicality for real-time deployment. In the CSI classification task, CSI-BERT2 consistently achieves the highest accuracy across various applications, including gesture recognition, person identification, fall detection, and crowd counting. Notably, our method shows strong robustness to discontinuous data caused by packet loss and to sampling rate mismatches between the training and testing sets, without the need to process and recover data or train a new model separately. Its accuracy drops by no more than 2%, in contrast to other models that suffer tens of percentage points in degradation. By demonstrating robust performance under packet loss and variable sampling rates, CSI-BERT2 eliminates the need for resource-intensive retraining in dynamic environments, making it viable for real-time applications such as emergency fall detection.
@@ III-A Overview
However, the original CSI-BERT does not fully leverage the advantages of MLM, resulting in subpar performance in downstream tasks. For example, the classification accuracy of CSI-BERT after fine-tuning does not surpass that of the vanilla ResNet18 [40]. This limitation can be attributed to the inadequacies of the embedding layer. To address this issue, CSI-BERT2 introduces an ARL-based spatial encoding module to capture the relationships and importance among subcarriers. Additionally, a MLP-based temporal encoding module is proposed to capture both the order and distance relationships among packets, effectively solving the temporal information loss problem inherent in conventional positional embedding modules, which capture distance relationships without considering order information. Moreover, to fully utilize the learned knowledge during pre-training, we propose a novel fine-tuning method named mask prediction model (MPM) for the CSI prediction task, wherein the last portion of the CSI sequence is randomly masked, and the model is trained to predict these masked values. Compared to conventional AR-based methods, our approach can complete the prediction in one step, thereby eliminating the error accumulation problem.
The model structure and working scenarios of CSI-BERT2 are illustrated in Fig. 1 and Fig. 2, respectively. We highlight the differences between CSI-BERT and CSI-BERT2 in Fig. 1 and Table I. We leverage the CSI sequence $c=[c_{1},c_{2},\ldots,c_{N}]$ and the corresponding timestamps $t=[t_{1},t_{2},\ldots,t_{N}]$ as input. Each $c_{i}$ represents a flattened vector of the CSI matrix at time $t_{i}$, which contains the amplitude and phase information of the subcarriers.
TABLE I: Comparison between CSI-BERT [11] and CSI-BERT2: Here PE represents the original positional embedding in Transformer, and SFT represents supervised fine-tuning. We utilize underline to highlight the differences between CSI-BERT2 and CSI-BERT.
CSI-BERT [11] CSI-BERT2 Model Structure Spatial Embedding MLP MLP + ARL Temporal Embedding PE PE + MLP Positional Embedding PE PE Backbone BERT BERT Discriminator DANN-based GAN-based Training Method Pre-training MLM MLM Classification Fine-tuning SFT SFT + Random Mask Prediction Fine-tuning $\times$ MPM Application Scenarios CSI Recovery $\checkmark$ $\checkmark$ CSI Classification Single Sampling Rate Various Sampling Rates CSI Prediction $\times$ $\checkmark$
@@ III-B Data Pre-processing
Due to packet loss, the received CSI series are always incomplete. Before inputting them to our model, we need to first identify where the packet loss happens and use a placeholder [PAD] to occupy the place of lost packets. We determine the occurrence of packet loss by analyzing the time gap between two successive timestamps. Given a sampling rate $f$, the ideal time gap between two successive timestamps should be $\Delta t=\frac{1}{f}$ seconds. Assume we receive a CSI series $c_{1},c_{2},\dotsc,c_{n}]$ with corresponding timestamps $t_{1},t_{2},\dotsc,t_{n}]$ with the unit of seconds. Then we can identify whether there is packet loss and how many packets were lost between $c_{i}$ and $c_{i+1}$. We calculate the lost packet amount according to:
@@ III-C1 Spatial and Temporal Embedding Layer
Compared to the spatial encoding, which primarily focuses on the inner information within each packet, the temporal encoding aims to provide more order and position information to the BERT backbone for better learning of the relationships among packets. However, we noted that the vanilla positional embedding in Transformers is insufficient. First, the arrival gap between successive packets is not uniform due to various factors such as environmental noise and movement. The positional embedding cannot capture these small differences. Most importantly, it has been shown to be weak in handling time series data [23], which can lead to temporal information loss.
@@ III-C2 Dual-branch Head Layer
For the sequence-level head, we use the same architecture as most Transformer-based methods [43, 44], which first incorporates a modified self-attention layer to combine the features in the time dimension and then uses a linear layer to generate the final output:
$x$ $\in\mathbb{R}^{N,d_{1}}\ ,$ (8)
$M_{attn}$ $=\text{Softmax}(\text{MLP}(x))\in\mathbb{R}_{+}^{N,d_{2}}\ ,$
$z$ $=\text{Flatten}(M_{attn}^{T}\cdot x)\in\mathbb{R}^{d_{1}\times d_{2}}\ ,$
$y$ $=\text{MLP}(z)\in\mathbb{R}^{o}\ ,$
where $d_{1}$ and $d_{2}$ represent the feature dimensions, $o$ denotes the output dimension, and $x$, $z$, and $y$ correspond to the input, latent feature, and output, respectively.
@@
Finally, considering the significance of real-time CSI classification and prediction models, we further analyze the time complexity of CSI-BERT2. The method primarily consists of self-attention and linear layers, so the complexity can be approximately expressed as $O(\alpha N^{2}d+\beta Nd^{2})$, where $N$ is the sequence length, $d$ represents the feature dimension, $\alpha$ is the number of self-attention layers, and $\beta$ is the number of linear layers. For simplification, we neglect the dimensional differences between the input, output, and hidden layers. Other computations, such as positional embedding encoding, dot products in ARL, and activation functions, can be absorbed by the two main terms. The self-attention complexity mainly arises from the BERT backbone, while the linear layers come from the BERT backbone, the embedding layer, and the output heads.
@@ III-D Unsupervised Pre-training Process
For CSI sequence, the MLM involves filling empty positions with [PAD] tokens, randomly masking tokens with [MASK] tokens, and training the model to recover the masked tokens. In BERT, the value of [MASK] token is fixed, but in CSI-BERT2, we replace the word embedding layer with a linear layer, which means the [MASK] token can take on different values affecting the output. To address this, we assign a random value to [MASK] tokens, sampled from a Gaussian distribution $MASK]_{i}^{(j)}\sim\text{N}(\mu_{i}^{(j)},\sigma_{i}^{(j)})$, where the mean and standard deviation are calculated from the standardization layer. This random approach makes it more challenging for the model to identify the positions of the [MASK] tokens, thereby encouraging a deeper understanding of the CSI sequence structure. For [PAD] token, even though we use a fixed value for [PAD], we replace them with [MASK] tokens before feeding the CSI sequence to the model. This is because there is no [PAD] in the CSI recovery phase. Even though the [PAD] can be ignored by the attention mechanism through the attention mask matrix, it still affects the output of the ARL, which works in the dimension of subcarrier. If [PAD] tokens only appear in the unsupervised training phase, it would cause a gap between inference and training, thereby influencing the model’s performance. As a result, we also disable the attention mask in this phase.
@@
$\underset{R}{\text{min}}\ \underset{D}{\text{max}}\ V(D,R)$ (9)
$=\underset{R}{\text{min}}\ \underset{D}{\text{max}}\ E_{c}[log(D(c))]+E_{c}[log(1-D(R(c)))]\ ,$
where $V$ is the target function, $D$ and $R$ represent the discriminator and recoverer, respectively, and $c$ represents the original CSI input. The discriminator tries to distinguish the original CSI input and the recovered CSI sequence output by the recoverer. The recoverer, in turn, learns to generate outputs that are indistinguishable from real CSI.
@@
[…]
@@
To reduce overfitting, we also modify the training data selection strategy. Unlike CSI-BERT, which uses a sliding window, CSI-BERT2 randomly samples CSI sequences from the entire dataset during training. The detailed pre-training algorithm is presented in Algorithm 1.
Algorithm 1 Unsupervised Pre-training Process
0: CSI dataset $\mathcal{D}=\{(c_{i},t_{i})\}_{i=1}^{N}$, pre-training epochs $E$, batch size $B$
0: Pre-trained CSI-BERT2 model $R$ (recoverer)
1: Initialize recoverer $R$ and discriminator $D$ with random weights
2: Preprocess data using Eq. (2) and (3) to handle packet loss
3:for epoch =1=1 to $E$ do
4:  for each batch of CSI sequences $\{(c^{(j)},t^{(j)})\}_{j=1}^{B}$ do
5:   Standardize each CSI sequence using Eq. (5)
6:   Randomly select masking ratio $r\sim\text{Uniform}(0.15,0.70)$
7:   Replace [PAD] tokens with [MASK] tokens
8:   Randomly mask $r$ proportion of non-[PAD] tokens with [MASK] tokens
9:   Assign random values to [MASK] tokens: $MASK]_{i}^{(j)}\sim\mathcal{N}(\mu_{i}^{(j)},\sigma_{i}^{(j)})$
10:   $\hat{c}\leftarrow R(c,t)$ {Forward pass through recoverer}
11:   Calculate recoverer loss $L_{R}$ using Eq. (10)
12:   Update $R$ by minimizing $L_{R}$ via backpropagation
13:   Calculate discriminator loss $L_{dis}$ using Eq. (11)
14:   Update $D$ by minimizing $L_{dis}$ via backpropagation
15:  end for
16:end for
17:return Pre-trained model $R$
@@ III-E1 CSI Prediction
In the inference phase of the prediction task, we first use [MASK] to replace all [PAD] tokens, and then we fix the last 20% of tokens in the input sequences as [MASK]. The model is then used to predict the last 20% of tokens. The detailed algorithm is similar to Algorithm 1, with the only difference being that the mask ratio $r$ is sampled from Uniform(0.15, 0.40), and the last $r$ tokens are replaced by [MASK].
@@ III-E2 CSI Classification
For CSI classification tasks, we adopt the approach commonly used for pre-trained models, where we freeze the bottom layers of CSI-BERT2 and only train the top layers for task-specific adaptation. Different from CSI-BERT, in this phase, we use [MASK] tokens to replace not only all [PAD] tokens but also some other input tokens randomly, to imitate the situation of different packet loss rates. This improves generalization to real-world scenarios where packet loss rates may vary during deployment. The detailed algorithm is presented in Algorithm 2.
@@
Algorithm 2 CSI Classification Fine-tuning
0: Pre-trained CSI-BERT2 model $R$, dataset $\mathcal{D}=\{(c_{i},t_{i},y_{i})\}_{i=1}^{N}$, fine-tuning epochs $E_{c}$, batch size $B$
0: Fine-tuned CSI classification model $F\circ R$
1: Freeze bottom layers of CSI-BERT2 $R$, only train top classification head $F$
2: Preprocess data using Eq. (2) and (3)
3:for epoch =1=1 to $E_{c}$ do
4:  for each batch $\{(c^{(j)},t^{(j)},y^{(j)})\}_{j=1}^{B}$ do
5:   Standardize each CSI sequence using Eq. (5)
6:   Randomly select masking ratio $r\sim\text{Uniform}(0.15,0.70)$
7:   Replace [PAD] tokens with [MASK] tokens
8:   Randomly mask $r$ proportion of non-[PAD] tokens with [MASK] tokens
9:   Assign random values to [MASK] tokens: $MASK]_{i}^{(j)}\sim\mathcal{N}(\mu_{i}^{(j)},\sigma_{i}^{(j)})$
10:   $\hat{y}\leftarrow F(R(c,t))$ {Classification output}
[…]
12:   Update classification head $F$ by minimizing $L$ via backpropagation
13:  end for
14:end for
15:return Fine-tuned classification model $F\circ R$